OpenStudy (jkbo):

Please help Solve: 15 decreased by twice a number is 21.  A.-3 B.3 C.18 D.-18

OpenStudy (mathstudent55):

Use the statement to get an equation. Then solve the equation. Let the number be x. 15 decreased by twice a number is 21. 15 - 2x = 21 Solve 15 - 2x = 21 for x.

OpenStudy (jkbo):

-3 @mathstudent55

OpenStudy (mathstudent55):

Correct. The answer is A.

OpenStudy (jkbo):

Can you help me with one more Solve: The sum of 7 times a number and 3 is 24. A.7/3 B.3 C.3/7 D.1/3 

OpenStudy (jkbo):

@mathstudent55

OpenStudy (mathstudent55):

Sure,

OpenStudy (mathstudent55):

Like with the previous problem, the first step is to choose a variable to represent the unknown quantity. Let x = the unknown number.

OpenStudy (mathstudent55):

Now you read the sentence and translate the words into an equation.

OpenStudy (mathstudent55):

a number = x 7 times a number = 7x the sum of 7 times a number and 3 = 7x + 3

OpenStudy (mathstudent55):

The sum of 7 times a number and 3 is 24. 7x + 3 = 24 Solve 7x + 3 = 24 for x

OpenStudy (jkbo):

@mathstudent55 3?

OpenStudy (mathstudent55):

Correct.

OpenStudy (jkbo):

@mathstudent55 one more thing please? 2/3 · 3/2 = 1 is an example of which property? Multiplicative Inverse  Reflexive  Inverse Operation  Identity of Multiplication 

OpenStudy (mathstudent55):

Reciprocals are also called multiplicative invereses. It's two numbers whose product is 1.

OpenStudy (jkbo):

ok so its A. @mathstudent55

OpenStudy (mathstudent55):

correct